Nomenclature with longtable

Brian Elmegaard April 7, 2000

Nomenclature

Roman Letters

f Residual value vector
J Jacobian Matrix
Δx Variable displacement vector
x Variable value vector

Subscripts

i  Variable number

0.1 Introduction

This file gives a very limited introduction to the use of longtable together with the nomencl package for writing nomenclatures.

It requires nomencl to be present on your system.

The way to go is: Install nomenltb.sty where latex can find it, and nomenltb.ist where makeindex can find that.

• Run latex. This generates the glossary input file foo.glo.

  • Run makeindex with makeindex -s nomenltb.ist -o foo.gls foo.glo

• Run latex again.